OpenStudy (anonymous):

simplify 12x + 36 over x^2 - 4x - 21 my answer: 12 over x - 7 x ≠ 7, x ≠ -3

OpenStudy (anonymous):

@anas2000

OpenStudy (anonymous):

please

OpenStudy (anonymous):

ill try

OpenStudy (anonymous):

12(x+3)/(x-7)(x+3) =12/(x-7) = 12(x+3)/(x+3)(x-7) =12/(x-7) x≠−3,x≠7

OpenStudy (anonymous):

wowww i was a negative sign off when i re-did it on my paper;....thanks